Accelerator Update for April 23, 2008 - April 25, 2008
Wednesday April 23
The day shift began with Operations monitoring the Recycler stash (201.4E10), with the Antiproton Source (Pbar) stacking (31.2E10), with the Recycler, the Main Injector, the Booster and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) waiting for the colliding beam, and with the NuMI and SciBooNE / MiniBooNE (BNB) experiments taking beam.
A Pbar expert began conducting target scans at 8:12 AM.
At 8:24 AM, Operations reported that two MI beam valves closed (BV520 & 523) and a P1 transfer line beam valve closed (BV701). Operators investigated and found the vacuum good and the permits good. Operators opened the valves, but they closed again after the first pulse of beam. Experts investigated.
Operations put Pbar into standby at 8:28 AM.
At 9:08 AM, Operations reported that the Recycler stash was gone. Two Recycler beam valves closed (BV610 & 601D). The valves closed while experts were replacing what they suspected was a bad control card for the MI and P1 beam valves, which had closed earlier. The Recycler valves got beam permit status from this crate. An expert reported that the 5-volt power supply for the crate was weak and needed replacing. Stacking, stashing, and NuMI beam was on hold until this work was completed around 10:38 AM.
CDF reported at 12:13 PM that they were troubleshooting their COT system and wanted two hours of quiet time.
At 12:21 PM, the TeV suffered a sector E4 quench due to a lead failure. There was no beam.
Operations put Pbar into standby at 1 PM.
Operations prepared D0 and CDF for an access at 1:18 PM.
A Recycler expert began some damper studies at 1:24 PM.
CDF came out of access at 2:26 PM.
At 3:20 PM, KRF3 tripped off. It reset.
At 3:24 PM, the TeV sector A4 cryo system reported it was not ready for ramping. This locked the system at 90 GeV. TECAR brought the ramp to flattop. The cryo system permit tripped off due to a valve closing (EVQH) while the TeV was ramping, which resulted in a JT valve (EVJT) opening to 55%. The Dewar (DT19) temperature rose to 4.5 Kelvin and drained to 20%. The Cryo Coordinator was informed and he began an investigation.
The Cryo Coordinator reported at 3:30 PM that the TeV sectors A4 and B1 cryo systems were stable.
Water experts reported at 3:58 PM that all the MI pond water and LCW pumps were running in their proper configuration.
Pbar resumed stacking at 4:06 PM.
Operations reported at 4:14 PM that the Recycler had stashed antiprotons twice during the day shift.
Operations put the TeV through a dry squeeze at 4:21 PM.
At 4:23 PM, MIRF2 tripped off on a bias power supply ground fault for the first of two times in four minutes. An expert investigated and had it back up about 20 minutes later.
At 4:53 PM, the NuMI beam permit tripped off due to MIRF2 tripping off. Operators investigated. Everything reset.
At 5:15 PM, a RF expert reported that there could be arcing on the MIRF1 or MIRF2 cavities in the tunnel. A RF expert will investigate further after Operations establishes the next store.
Operations put Pbar into standby mode at 5:31 PM.
A RF expert reported at 6:10 PM that the MIRF1 cavity was the one that would have to be inspected.
Operations put the TeV into shot setup at 7:44 PM.
At 7:46 PM, KRF6 tripped off. It reset.
At 8:02 PM, KRF2 tripped off. It reset.
Operations began preparing MI/Recycler at 9:43 PM for an access.
At 9:55 PM, three MI beam valves closed (BV701, BV706, and BVF11) for no reason. Their permits were good. A controls expert was called in.
Operations established store 6076 at 10:04 PM with an initial luminosity of 128.02E30.
At 10:15 PM, a MI expert reported that he was working on the MI-52 LCW PLC problem.
RF experts accessed MI at 10:32 PM.
A controls expert reported at 11:38 PM that he had replaced a control card, which should resolve the MI beam valve closings.
RF experts reported at 11:43 PM that they found the top of the MIRF1 cavity pitted due to arcing. They began reforming the flange, replacing the gasket, and the station’s Power Amplifier (PA).
Operations reported at midnight that the Recycler had stashed twice during the evening shift.
Thursday April 24
The midnight shift began with Operations monitoring store 6076, with MI/Recycler in access, with Operations monitoring the Pbar stack (52.6E10), and with BNB taking beam.
At 12:09 AM, the TeV quenched. Operators investigated.
Operations reported at 12:31 AM that a TeV sector D0 quadrupole power supply (D0Q6) was the cause of the TeV quench. Operations called in a power supply expert.
Operations had CDF prepared for an access by 1:13 AM.
At 2:20 AM, KRF1 tripped off. It reset.
MI came out of access at 2:32 AM.
A TeV power supply expert reported at 3:54 AM that he put a new PLC chassis in D0Q5 power supply and the put the old Q5 PLC chassis into D0Q6.
At 4:29 AM, LRF4 tripped off. It reset.
Operations reported at 4:31 AM that they had begun the TeV turn on.
Operations put the TeV through a dry squeeze at 5:06 AM.
At 6:03 AM, Operations reported that the last two stashes had a very poor efficiency. Operators believed that the problem was due to a correction element not ramping and a Pbar expert concurred with that theory. Unfortunately, that element hadn’t been found yet and so the investigation continued.
At 6:47 AM, LRF2 tripped off and would not reset. Operations called in a Linac expert.
Operations put the TeV through a wet squeeze at 6:51 AM.
Operations put Pbar into standby at 7:12 AM.
Operations reported that LRF2 was operating normally by 7:57 AM.
Operations reported at 8 AM that the Recycler had stashed twice during the midnight shift.
At 8:08 AM, LRF2 tripped off for the first of two times in about an hour. It reset.
Cryo system experts reported at 10:14 AM that they were beginning wet engine maintenance at TeV sectors F1 and F2.
At 11:44 AM, KRF7 tripped off. It reset.
At 11:54 AM, KRF2 tripped off. It reset.
Operations put the TeV through a dry squeeze at 12:15 PM.
At 12:42 PM, Operations reported that two MI CAMAC crates (10 & 17) stopped reporting. Due to the loss of data from these crates, the Recycler stash had lost some antiprotons, but less than 1E10. Operators investigated and contacted experts.
Operations put Pbar into standby at 12:45 PM.
Operators reported at 1 PM that they found MI CAMAC crate 10 off and with a strong smell of burned electronics. Controls experts replaced a power supply and a bad card. There was nothing wrong with CAMAC crate 17.
Operations established beam through MI at 1:12 PM.
A Booster expert requested at 1:18 PM that Operations try to send NuMI eight turns of beam. Operators succeeded in increasing the beam intensity to NuMI. Two MI experts began tuning on the their accelerator.
At 1:41 PM, KRF2 tripped off. It reset.
Operations put the TeV into shot setup at 1:57 PM.
At 3:04 PM, KRF7 tripped off. It reset.
At 3:40 PM, A Booster expert reported a timing problem with a Booster kicker (MKS06). He said that beam would be off for a few minutes while an expert changed a timing module.
Operations reported at 3:56 PM that the Recycler had stashed three times during the day shift.
Operations established store 6080 at 4:02 PM with an initial luminosity or 144.1E30.
At 4:41 PM, KRF6 tripped off for the first to two times in about an hour. It reset.
At 5:05 PM, KRF7 tripped off for the first of two times in about 90 minutes. It reset.
At 6:21 PM, KRF2 tripped off. It reset.
At 7:23 PM, KRF3 tripped off. It reset.
At 11:21 PM, LRF2 tripped off. It reset.
Operations reported at midnight that the Recycler had stashed six times during the evening shift.
Friday April 25
The midnight shift began with Operations monitoring store 6080 (54.41E30) and the stash (181.8E10), with Pbar stacking (21.1E10), and with NuMI and BNB taking beam.
At 12:37 AM, a TeV air compressor (T:A1AIRP) alarmed. It reset. Operators investigated and found its tank blowdown valve had stuck open.
At 1:51 AM, CDF reported that they had a failing calorimeter power supply and would need about a two-hour access to replace it. Operations informed the Run Coordinator and he approved their request.
At 2:14 AM, Operations reported another round of poor stashing efficiencies. Operators had good success with some adjustments.
Operations terminated store 6080 at 4:04 AM.
Operators prepared CDF for access at 4:22 AM.
CDF came out of access at 6:01 AM.
Operations put the TeV through a dry squeeze at 6:19 AM.
At 6:33 AM, the NuMI beam permit tripped off when a correction element (E:V100) power supply tripped off. Everything reset.
Operations put the TeV into shot setup at 6:43 AM.
Operations reported at 7:55 AM that the Recycler had stashed three times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.
Comments and Suggestions
|last modified 04/28/2008 email Fermilab|
|Security, Privacy, Legal|